Kyky232:

Triangle ABC has vertices located at A( 0, 2), B (2, 5), and C (−1, 7). Part A: Find the length of each side of the triangle. Show your work. Part B: Find the slope of each side of the triangle. Show your work. Part C: Classify the triangle. Explain your reasoning. am confused plz help

Kyky232:

plz help

Kyky232:

@AZ

simsharrison:



simsharrison:

thats what it looks like when you put it on a graph, the one thats NOT on the graph is not drawn to scale. I just did that one to show what it looks like and to show the angles.

simsharrison:

Idk, how to do A and B But for C, which triangle classification is less than 90*?

jhonyy9:

to find the length of sides just use the distance formula - hope you know it @supie

simsharrison:

@jhonyy9 to find the slope, would we use rise/run?

supie:

The distance formula as in \(\large{d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}}\) where d=distance (x1, x2) the x coordinates (y2, y1) the y coordinates @jhonyy9 ?

simsharrison:

@Kyky232

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie The distance formula as in \(\large{d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}}\) where d=distance (x1, x2) the x coordinates (y2, y1) the y coordinates @jhonyy9 ? \(\color{#0cbb34}{\text{End of Quote}}\) yes @supie

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @supie The distance formula as in \(\large{d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}}\) where d=distance (x1, x2) the x coordinates (y2, y1) the y coordinates @jhonyy9 ? \(\color{#0cbb34}{\text{End of Quote}}\) yes @supie \(\color{#0cbb34}{\text{End of Quote}}\) good job

supie:

Thanks